A semiconductor manufacturer produces controllers used in automobile engine applications. The customer required that the process fallout or fraction defective at a critical manufacturing step not exceed 0.05 and that the manufacturer demonstrate process capacity at this level of quality using α = 0.05. The semiconductor manufacturer takes a random sample of 200 devices and finds that four of them are defective. Can tha manufaturer demonstrate process capability for the customer?
1
Expert's answer
2020-05-14T16:08:09-0400

 We need to test "H_0:p=0.05" against "H_1:p<0.05." The test statistic is 


"z={X-np\over \sqrt{np(1-p)}}"

Given "n=200, X=4"


"z={4-200(0.05)\over \sqrt{200(0.05)(1-0.05)}}\approx-1.946657"


The critical value is "z_c=z_{1-\alpha}=-1.645." Rhe rejection region is "z<-1.645."

Since "z=-1.946657<-1.645," we reject "H_0," and we can conclude that there is enough evidence to support the fact that the process fallout does not exceed.

Note that p-value "P(Z\leq-1.946657)=0.025788."  Since p − value "=0.025788<0.05=\alpha," we reject "H_0," and we can conclude that there is enough evidence to support the fact that the process fallout does not exceed.